How is calculating sales tax similar to calculating a discount? Explain and give examples.
Answer #1

normaly, if sales price,percentage of discount/tax is given seperately we can apply the given percentage to sales price and can extract the discount/tax amount.for example

1)if selling price of a bag is $150 and discount rate is 45%,then to find discounted selling price,first convert the given percentage to decimel and multiply it with sales price.mathematicallly 45%=45/100=.45,hence discounted selling price=sale price-discount=sales price-sales pricexdiscount =150-150*.45=150-67.5=$82.5

also,if sales tax rate is 15%,and sales price is 150(inclusive of tax),the to extract the tax amount from the sales price.we need to use the formula,tax=sales price*rate o tax(100+rate of tax),in the given case it is =150x15/115=19.57

so sales price excluding tax=150-19.57=$130.43.

if sales price given exclusive of tax rate say 150 and tax rate is 20%,then we can find sales price inclusive of tax=sales price(excluding tax)+tax rate*sales price=150+20%x150=150+30=180.